OpenStudy (swissgirl):

Which of these sets have the inductive property: 1) {20,21,22,23,...} 2){1,2,4,5,6,7,...} 3) {\( x \in N : x^2 \le 1000 \)}

OpenStudy (dominusscholae):

1). It has a general pattern that allows us to predict the next number in the set.

OpenStudy (swissgirl):

what abt #3?

OpenStudy (dominusscholae):

Wait....I'm sorry. I think I got this confused with induction....sorry I don't know :(.

OpenStudy (swissgirl):

nooo. A set S of natural numbers is inductive iff it has the property that whenever \( n \in S \) then \( n+1 \in S \)

OpenStudy (swissgirl):

@SmoothMath can u help me?

OpenStudy (swissgirl):

I know that #1 is but there seems to be another set by the way the question was phrased/ So u think the #3 is an option too?

OpenStudy (anonymous):

Swissgirl, let's look at #3 and ask some questions. Is 31 a members of that set?

OpenStudy (swissgirl):

nope

OpenStudy (anonymous):

orly? Tell me why.

OpenStudy (swissgirl):

since there is no natural number,x that when u square it will give you 31

OpenStudy (anonymous):

You are misreading that set.

OpenStudy (swissgirl):

ohhhhh yaaaaaa ok i did lol

OpenStudy (anonymous):

{x∈N:x^2≤1000} means "x in N such that x^2 is less than or equal to 1000" That means that x is in the set if it satisfies those conditions. It has to be in N, that is, a natural number. And it has to square to be less than or equal to 1000.

OpenStudy (swissgirl):

yaaaaa ok so then its inductive

OpenStudy (anonymous):

D':

OpenStudy (swissgirl):

but 31 isnt part of the set

OpenStudy (swissgirl):

since 31^2=961

OpenStudy (anonymous):

31 isn't in the set, eh?

OpenStudy (swissgirl):

again i misread even what u wrote lol

OpenStudy (swissgirl):

so then its not inductive since it doesnt go on forever

OpenStudy (anonymous):

Swissgirl, there are two things that we need to understand here before we can answer this question, and I don't think you understand either of them. First we have to understand what it means for a a set to be inductive. And we need to understand what the set is.

OpenStudy (anonymous):

Take it slow and understand both of those things completely before you try to jump to an answer.

OpenStudy (anonymous):

Your understanding of inductive is actually pretty good. "It goes on forever" pretty much works. If there is a last element, then it is not inductive.

OpenStudy (swissgirl):

ya since if n is the last number then there wont be an n+1

OpenStudy (anonymous):

Good good.

OpenStudy (swissgirl):

so then its not inductive

OpenStudy (anonymous):

Explain to me your understanding of the set.

OpenStudy (swissgirl):

lets call this set S which is composed of natural numbers. Whenever \( n \in S \) then \(n+1 \in S \)

OpenStudy (anonymous):

Oh, I meant your understanding of what set 3 is.

OpenStudy (swissgirl):

ohhhhh lol

OpenStudy (swissgirl):

ok so the set S includes all natural numbers that when squared is less or equal to 1000

OpenStudy (swissgirl):

is less than *

OpenStudy (swissgirl):

so let n be an element of this set S. Let n=31 so 31^2=961 now n+1=32 and 32^2=1024> since n+1 isnt an element of this set so this set isnt inductive

OpenStudy (anonymous):

Very good =)

OpenStudy (swissgirl):

YAAYYYYYYYYYYY. Thanks

OpenStudy (anonymous):

So for any set, if n is in the set, but n+1 is not, then that set is not inductive.

OpenStudy (swissgirl):

yuppp gotchhaaa

OpenStudy (anonymous):

Which means that if some number is in the set, then every number after it has to be too. So as you said, it has to go on forever, at least from the starting point that you establish.

OpenStudy (swissgirl):

Thanks :)

OpenStudy (anonymous):

Mhmm =) Question. Is it the notation that they used for that third set that confused you? Can I help you out with how to read that kind of notation?

OpenStudy (swissgirl):

umm well kinda. I do have difficulties reading it plus i was chatting at the same time to be honest

OpenStudy (anonymous):

Okay, I understand. Just take your time and try to wrap your head around what the notation means. The : in the notation means "such that" and what comes after it should be thought of as a criteria for what can be included in the set.

OpenStudy (swissgirl):

okkkkkkkkkk gotcha thanks

OpenStudy (anonymous):

Yup =)